# Continuous operators from spaces of Lipschitz functions

### v1

@@ Abstract.
We study the existence of continuous (linear) operators from the Banach spaces $\operatorname{Lip}_{0}(M)$ of Lipschitz functions on infinite metric spaces $M$ vanishing at a distinguished point and from their predual spaces $\mathcal{F}(M)$ onto certain Banach spaces, including $C(K)$-spaces and the spaces $c_{0}$ and $\ell_{1}$. For pairs of spaces $\operatorname{Lip}_{0}(M)$ and $C(K)$ we prove that if they are endowed with topologies weaker than the norm topology, then usually no continuous (linear or not) surjection exists between those spaces. We show that, given a Banach space $E$, there exists a continuous operator from a Lipschitz-free space $\mathcal{F}(M)$ onto $E$ if and only if $\mathcal{F}(M)$ contains a subset homeomorphic to $E$ if and only if $d(M)\geq d(E)$. We obtain a new characterization of the Schur property for spaces $\mathcal{F}(M)$: a space $\mathcal{F}(M)$ has the Schur property if and only if for every discrete metric space $N$ with cardinality $d(M)$ the spaces $\mathcal{F}(M)$ and $\mathcal{F}(N)$ are weakly sequentially homeomorphic. It is also showed that if a metric space $M$ contains a bilipschitz copy of the unit sphere $S_{c_{0}}$ of the space $c_{0}$, then $\operatorname{Lip}_{0}(M)$ admits a continuous operator onto $\ell_{1}$ and hence onto $c_{0}$. We provide several conditions for a space $M$ implying that $\operatorname{Lip}_{0}(M)$ is not a Grothendieck space.
@@ 1. Introduction
To deal with this problem, in Section 3 we compute the density characters of spaces of Lipschitz functions endowed with various topologies. This is also the starting point for obtaining the characterization of homeomorphism classes of spaces $\operatorname{Lip}_{0}(M)$, presented in Corollary 3.5 and asserting that, given two metric spaces $M$ and $N$, the Banach spaces $\operatorname{Lip}_{0}(M)$ and $\operatorname{Lip}_{0}(N)$ are homeomorphic if and only if $2^{d(M)}=2^{d(N)}$, where $d(X)$ denotes the density of a topological space $X$ (see Section 2 for relevant definitions and notations).
The density of spaces $\operatorname{Lip}_{0}(M)$ endowed with topologies significantly weaker than the norm topology or even the weak topology are studied in Theorem 3.8 and its several corollaries, in particular Theorem 3.13. We show, among others, that if for a given metric space $M$ the space $\operatorname{Lip}_{0}(M)$ is endowed with the topology $\tau$ contained between the pointwise topology and the compact-open topology, then for the densities we have $d(C(M),\tau)=d(\operatorname{Lip}_{0}(M),\tau)\leq d(M)$. This provides a metric extension for Noble’s classical result [52], cf. also [33].
@@
In Theorem 4.5 we consider the reverse situation, which also appears to be negative. Namely, we show that for every infinite Tychonoff space $X$ and infinite metric space $N$ for which the inequality $w(X)<2^{d(N)}$ holds, there is no continuous (linear or not) surjection from $(C(X),\tau)$ onto $(\operatorname{Lip}_{0}(N),\tau^{\prime})$, where $\tau$ is a topology contained between the pointwise topology and the compact-open topology on $C(X)$ and a topology $\tau^{\prime}$ contains the weak topology of the Banach space $\operatorname{Lip}_{0}(N)$. The assumption $w(X)<2^{d(N)}$ cannot be removed nor even relaxed to $w(X)\leq 2^{d(N)}$, which is justified by the well-known isomorphisms $C(\beta\mathbb{N})\simeq\ell_{\infty}\simeq\operatorname{Lip}_{0}([0,1])$, where $\beta\mathbb{N}$ is the Čech–Stone compactification of the set $\mathbb{N}$ of natural numbers. On the other hand, it should be pointed out, that according to Cúth, Doucha, and Wojtaszczyk, see [19, proof of Theorem 4.1], there is no continuous linear surjection $T\colon C(K)\rightarrow\operatorname{Lip}_{0}([0,1]^{2})$ for any compact space $K$. This in particular yields that for any metric space $M$ containing the square $[0,1]^{2}$ there is no continuous linear surjection $T\colon C(K)\rightarrow\operatorname{Lip}_{0}(M)$ for any compact space $K$.
In Subsection 4.1 we study the existence of continuous mappings from or onto Lipschitz-free Banach spaces. In Theorem 4.6 we show that for an infinite metric space $M$ and a Banach space $E$ there exists a continuous (linear) surjection from the Lipschitz-free space $\mathcal{F}(M)$ onto $E$ if and only if $\mathcal{F}(M)$ contains a subset homeomorphic to $E$ if and only if $d(M)\geq d(E)$. Further, motivated by a recently studied open problem due to Krupski [46] (see also [47]), asking whether the spaces $C_{p}(K)$ and $C(L)_{w}$ can be homeomorphic for infinite compact spaces $K$ and $L$, where, for a Tychonoff space $X$ and a Banach space $E$, $C_{p}(X)$ denotes the space $C(X)$ endowed with the pointwise topology and $E_{w}$ denotes $E$ endowed with the weak topology, we discuss in Theorem 4.8 possible natural conditions put on two metric spaces $M$ and $N$ under which there is no homeomorphism between spaces $C_{p}(M)$ and $\mathcal{F}(N)_{w}$.
@@
We provide a new characterization of the Schur property of Lipschitz-free spaces, more in the spirit of the title of our paper. Namely, in Theorem 4.15 we prove that for an infinite metric space $M$ the space $\mathcal{F}(M)$ has the Schur property if and only if for every discrete metric space $N$ with cardinality $d(M)$ the spaces $\mathcal{F}(M)_{w}$ and $\mathcal{F}(N)_{w}$ are sequentially homeomorphic. The main tool used for proving this result is Theorem 4.10 stating that a Banach space $E$ has the Schur property if and only if the spaces $E$ and $E_{w}$ are sequentially homeomorphic. We also observe in Theorem 4.13 that, for every two Banach spaces $E$ and $F$ both having the Schur property, the spaces $E_{w}$ and $F_{w}$ are sequentially homeomorphic if and only if $d(E)=d(F)$.
Section 5 is devoted to the second main problem of this paper, motivated by the observation that for every infinite metric space $M$, the space $\operatorname{Lip}_{0}(M)$ contains a complemented copy of the Banach space $\ell_{\infty}(d(M))$ (by Hájek–Novotný’s result [35], see also Theorem 3.2), hence there is always a continuous linear surjection $T\colon\operatorname{Lip}_{0}(M)\rightarrow\ell_{2}(2^{d(M)})$ and thus a continuous linear surjection $S\colon\operatorname{Lip}_{0}(M)\rightarrow\ell_{2}$, see Corollary 5.1.
@@
Using Corollary 5.7 and the theorem of Dalet [23] stating that the space $\operatorname{Lip}_{0}(\ell_{1})$ contains a complemented copy of $\ell_{1}$ (see Proposition 5.9), we provide in Theorem 5.10 an extensive list of examples of Banach spaces $E$ for which the spaces $\operatorname{Lip}_{0}(E)$ admit continuous operators onto $\ell_{1}$. To this list belong, e.g., the following Banach spaces:
  • predual spaces or dual spaces of arbitrary order of C*-algebras, in particular, all $C(K)$-spaces, $L_{1}(\mu)$-spaces, and $L_{\infty}(\mu)$-spaces,
@@
It seems that apart from $\operatorname{Lip}_{0}([0,1])\simeq\ell_{\infty}$ there is no known example of a Banach space $\operatorname{Lip}_{0}(M)$ which is a Grothendieck space. Recall that for the class of $C(K)$-spaces we have the following result, due to Cembranos [16] and Räbiger [55]: for every compact space $K$, the Banach space $C(K)$ is Grothendieck if and only if $C(K)$ does not contain any complemented copy of $c_{0}$. This characterization cannot work for spaces $\operatorname{Lip}_{0}(M)$, since for each metric space $M$ the space $\operatorname{Lip}_{0}(M)$ is isometrically isomorphic to the dual Banach space $\mathcal{F}(M)^{*}$, hence it does not contain any complemented copy of $c_{0}$ (see e.g. [21, Theorem 2.4.15]), even though, by Theorem 3.2, $\operatorname{Lip}_{0}(M)$ always contains some copy of $c_{0}$. It follows that to decide that a given space $\operatorname{Lip}_{0}(M)$ is not Grothendieck we have to look for continuous linear surjections from $\operatorname{Lip}_{0}(M)$ onto $c_{0}$ or prove that the dual space $\operatorname{Lip}_{0}(M)^{*}$ is not weakly sequentially complete; see the discussion at the beginning of Section 5.1 for more details.
In Subsection 5.1 we provide a number of conditions for metric spaces $M$ implying that their spaces $\operatorname{Lip}_{0}(M)$ are not Grothendieck. For example, since the Grothendieck property is preserved by continuous linear surjections and the space $\operatorname{Lip}_{0}(E)$ contains a complemented copy of the dual $E^{*}$ for each Banach space $E$, the space $\operatorname{Lip}_{0}(E)$ is not Grothendieck provided that $E^{*}$ is not Grothendieck. This observation yields a class of Banach spaces $E$ for which $\operatorname{Lip}_{0}(E)$ is not Grothendieck, e.g. all non-reflexive spaces with separable duals belong to this class, see Proposition 5.13.
@@ 2. Preliminaries
For a Banach space $E$, the closed unit ball of $E$ will be denoted by $B_{E}$. $E^{*}$ denotes the dual Banach space of $E$ and $E_{w}$ denotes $E$ endowed with the weak topology, i.e. the topology $w=\sigma(E,E^{*})$. The weak* topology of $E^{*}$ will be denoted by $w^{*}$, i.e. $w^{*}=\sigma(E^{*},E)$.
@@
The Banach spaces $c_{0}$ and $L_{p}(\Omega,\Sigma,\mu)$ ($1\leq p\leq\infty$), in particular $\ell_{1}(\Gamma)$ and $\ell_{\infty}(\Gamma)$ for a set $\Gamma$ and $\ell_{\infty}^{n}$ for an integer $n\in\mathbb{N}$, are defined in the standard way.
For a topological space $X$, by $w(x)$ we denote the weight of $X$ (i.e. the minimal cardinality of a base of the topology of $X$), by $d(X)$ the density of $X$ (i.e. the minimal cardinality of a dense subset of $X$), by $d_{seq}(X)$ the sequential density of $X$ (i.e. the minimal cardinality of a sequentially dense subset of $X$, that is, such a subset $A\subseteq X$ that for every $x\in X$ there exists a sequence $(x_{n})_{n=0}^{\infty}$ in $A$ converging to $x$), by $c(X)$ the cellularity of $X$ (i.e. the supremum of cardinalities of families of pairwise disjoint non-empty open subsets of $X$), and by $ww(X)$ the weak weight of $X$ (i.e. the minimal weight $w(Y)$ of a Tychonoff space $Y$ such that there exists a continuous bijection $\varphi\colon X\rightarrow Y$). We always have $c(X)\leq d(X)\leq w(X)$, $d(X)\leq d_{seq}(X)$, and $w(X)\geq ww(X)$. If $X$ is metric, then we have $c(X)=d(X)=d_{seq}(X)=w(X)\geq ww(X)$, and the cellularity is attained, that is, there exists a collection of pairwise disjoint non-empty open subsets of $X$ of cardinality $c(X)$ (see [38, Theorem 8.1]).
@@ 3. Density of spaces of Lipschitz functions
In this paper we will frequently use the fact that for every infinite metric space $M$ the spaces $\operatorname{Lip}(M)$ and $\operatorname{Lip}_{0}(M)$ contain isomorphic copies of the Banach space $\ell_{\infty}(d(M))$, where $d(M)$ denotes the density of $M$ (see Theorems 3.1 and 3.2). These results can be easily deduced from the proof of Hájek and Novotný’s [35, Proposition 3] when combined with Rosenthal’s [59, Corollary 1.2]. Since our proof is somewhat more direct, we add it for the convenience of the reader and to keep this paper more self-contained. The main difference between our proof and the one in [35] is that we use the cellularity of $M$ to obtain directly a suitable family of disjoint open balls, which makes the argument simpler.
@@
Since $M$ is a metric space, its cellularity $c(M)$ is equal to the density $d(M)$, and so we may pick a family of pairwise disjoint (open) balls $(B(y_{\alpha},r_{\alpha}))_{\alpha<d(M)}$ of cardinality $d(M)$ such that $r_{\alpha}\leq 1$ for every $\alpha<d(M)$ (see [38, Theorem 8.1]). For each $\alpha<d(M)$ we set
@@
where $e_{\alpha}$ denotes the $\alpha$-th standard basic vector of $\ell_{\infty}(d(M))$, we may use Rosenthal’s [34, Theorem 7.10] to conclude the existence of a set $\Gamma$ with cardinality $d(M)$ such that $\ell_{\infty}(\Gamma)$ is isomorphic to a subspace of $\operatorname{Lip}(M)$. This finishes the proof. ∎
Following [64, Definition 2.17], for a complete metric space $M$ with the metric $\rho$ and the base point $e$, the space $M^{\dagger}$ is defined as $M\setminus\{e\}$ and equipped with the metric
@@
for $x,y\in M\setminus\{e\}$. By [64, Theorem 2.20], the Banach spaces $\operatorname{Lip}_{0}(M)$ and $\operatorname{Lip}(M^{\dagger})$ are isomorphic—we will use this result together with Theorem 3.1 to deduce the following theorem of Hájek and Novotný [35], which will constitute one of our main tools in what follows.
@@
The latter inequality yields that $\rho(x_{n},x)\rightarrow 0$ as $\rho^{\dagger}(x_{n},x)\to 0$. It follows that the inverse mapping $f^{-1}$ is also sequentially continuous and hence continuous. Consequently, $f$ is a homeomorphism between $M\setminus\{e\}$ and $M^{\dagger}$, and so $d(M)=d(M^{\dagger})$. By aforementioned [64, Theorem 2.20], the Banach spaces $\operatorname{Lip}_{0}(M)$ and $\operatorname{Lip}(M^{\dagger})$ are isomorphic and hence the claim follows from Theorem 3.1. ∎
As a direct consequence of Theorem 3.2 we obtain Hájek and Novotný’s ([35, Proposition 3]).
@@
Since by Theorem 3.2 the dual space $\operatorname{Lip}_{0}(M)$ of $\mathcal{F}(M)$ contains a copy of $\ell_{\infty}(d(M))$, the result follows from [59, Corollary 1.2]. ∎
@@
We already know that $d(\operatorname{Lip}_{0}(M))\geq 2^{d(M)}$ and $d(\operatorname{Lip}(M))\geq 2^{d(M)}$. The other direction follows from the fact that there are at most $2^{d(M)}$ many continuous real-valued functions on $M$, see e.g. [38, Theorem 10.1]. ∎
The next corollary immediately follows from the classical Toruńczyk theorem ([63]) stating that two infinite-dimensional Banach spaces are homeomorphic if and only if they have the same density.
@@
Of course, if $d(M)=d(N)$, then $2^{d(M)}=2^{d(N)}$, however the converse implication in general does not hold, as there are models of set theory (like original Cohen’s model for violating the Continuum Hypothesis, see [48, Chapter VII]) in which the equality $2^{\aleph_{0}}=2^{\aleph_{1}}$ is satisfied.
@@
Assume that $E$ is a closed linear subspace of $\operatorname{Lip}_{0}(M)_{p}$ isomorphic to $(c_{0})_{p}$. Since the closed unit ball $B$ of $\operatorname{Lip}_{0}(M)$ is compact in $\operatorname{Lip}_{0}(M)_{p}$, for the sequence $(B_{n})_{n=0}^{\infty}$ of compact sets, where $B_{n}=nB$ for $n\in\mathbb{N}$, we obtain the compact cover $(B_{n}\cap E)_{n=0}^{\infty}$ of $E$. On the other hand, the space $(c_{0})_{p}$ is isomorphic to the space $C_{p}(\mathbb{N}^{\#})$, where $\mathbb{N}^{\#}$ is a one-point compactification of the discrete space $\mathbb{N}$ of natural numbers. It follows that $C_{p}(\mathbb{N}^{\#})$ is covered by a sequence of compact sets. This however yields a contradiction by [5, Theorem I.2.1]. ∎
@@
We will now compute densities of spaces of Lipschitz functions endowed with various topologies weaker than the compact-open topology. The next theorem is an extension of Noble’s [52, Theorem 1].
@@
Since $M$ is metric, $d(M)=w(M)\geq ww(M)$. By [52, Theorem 1], $ww(M)=d(C(M),\tau)$.
@@
We have $G^{U,V}_{n}\subseteq G^{U,V}_{n+1}\subseteq G^{U,V}$. Also, if $G^{U,V}_{n}\neq\emptyset$, then $\rho(x,G^{U,V}_{n})\geq 1/n$ for every $x\in H^{U,V}$.

### v3

@@ Abstract.
We study the existence of continuous (linear) operators from the Banach spaces $\operatorname{Lip}_{0}(M)$ of Lipschitz functions on infinite metric spaces $M$ vanishing at a distinguished point and from their predual spaces $\mathcal{F}(M)$ onto certain Banach spaces, including $C(K)$-spaces and the spaces $c_{0}$ and $\ell_{1}$. For pairs of spaces $\operatorname{Lip}_{0}(M)$ and $C(K)$ we prove that if they are endowed with topologies weaker than the norm topology, then usually no continuous (linear or not) surjection exists between those spaces. It is also showed that if a metric space $M$ contains a bilipschitz copy of the unit sphere $S_{c_{0}}$ of the space $c_{0}$, then $\operatorname{Lip}_{0}(M)$ admits a continuous operator onto $\ell_{1}$ and hence onto $c_{0}$. Using this, we provide several conditions for a space $M$ implying that $\operatorname{Lip}_{0}(M)$ is not a Grothendieck space. Finally, we obtain a new characterization of the Schur property for Lipschitz-free spaces: a space $\mathcal{F}(M)$ has the Schur property if and only if for every complete discrete metric space $N$ with cardinality $d(M)$ the spaces $\mathcal{F}(M)$ and $\mathcal{F}(N)$ are weakly sequentially homeomorphic.
@@ 1. Introduction
To deal with this problem, in Section 3 we compute the density characters of spaces of Lipschitz functions endowed with various topologies. For the norm topology, using Toruńczyk’s classical theorem ([65]), we observe in Corollary 3.5 that, given two metric spaces $M$ and $N$, the Banach spaces $\operatorname{Lip}_{0}(M)$ and $\operatorname{Lip}_{0}(N)$ are homeomorphic if and only if $2^{d(M)}=2^{d(N)}$, where $d(X)$ denotes the density of a topological space $X$ (see Section 2 for relevant definitions and notations). The density of spaces $\operatorname{Lip}_{0}(M)$ endowed with topologies significantly weaker than the norm topology or even the weak topology are studied in Theorem 3.8 and its several corollaries, in particular Theorem 3.13. We show, among others, that if for a given metric space $M$ the space $\operatorname{Lip}_{0}(M)$ is endowed with the topology $\tau$ contained between the pointwise topology and the compact-open topology, then for the densities we have $d(C(M),\tau)=d(\operatorname{Lip}_{0}(M),\tau)\leq d(M)$. This provides a metric extension for Noble’s theorem [55], cf. also [35], as well as relates to results presented in the recent paper [14] which concern the density character of the weak* topology on spaces of Lipschitz functions.
@@
In Theorem 4.5 we consider the reverse situation, which also appears to be negative. Namely, we show that for every infinite Tychonoff space $X$ and infinite metric space $N$ for which the inequality $w(X)<2^{d(N)}$ holds, there is no continuous (linear or not) surjection from $(C(X),\tau)$ onto $(\operatorname{Lip}_{0}(N),\tau^{\prime})$, where $\tau$ is a topology contained between the pointwise topology and the compact-open topology on $C(X)$ and a topology $\tau^{\prime}$ contains the weak topology of the Banach space $\operatorname{Lip}_{0}(N)$. The assumption $w(X)<2^{d(N)}$ cannot be removed nor even relaxed to $w(X)\leq 2^{d(N)}$, which is justified by the well-known isomorphisms $C(\beta\mathbb{N})\simeq\ell_{\infty}\simeq\operatorname{Lip}_{0}([0,1])$, where $\beta\mathbb{N}$ is the Čech–Stone compactification of the set $\mathbb{N}$ of natural numbers. On the other hand, it should be pointed out, that according to Cúth, Doucha, and Wojtaszczyk, see [21, proof of Theorem 4.1], there is no continuous linear surjection $T\colon C(K)\rightarrow\operatorname{Lip}_{0}([0,1]^{2})$ for any compact space $K$. This in particular yields that for any metric space $M$ containing the square $[0,1]^{2}$ there is no continuous linear surjection $T\colon C(K)\rightarrow\operatorname{Lip}_{0}(M)$ for any compact space $K$.
In Subsection 4.1 we study the existence of continuous mappings from or onto Lipschitz-free Banach spaces. In Proposition 4.6 we observe that, again by virtue of the Toruńczyk theorem, for an infinite metric space $M$ and a Banach space $E$ there exists a continuous (linear) surjection from the Lipschitz-free space $\mathcal{F}(M)$ onto $E$ if and only if $\mathcal{F}(M)$ contains a subset homeomorphic to $E$ if and only if $d(M)\geq d(E)$. Further, motivated by a recently studied open problem due to Krupski [49] (see also [50]), asking whether the spaces $C_{p}(K)$ and $C(L)_{w}$ can be homeomorphic for infinite compact spaces $K$ and $L$, where, for a Tychonoff space $X$ and a Banach space $E$, $C_{p}(X)$ denotes the space $C(X)$ endowed with the pointwise topology and $E_{w}$ denotes $E$ endowed with the weak topology, we discuss in Theorem 4.8 possible natural conditions put on two metric spaces $M$ and $N$ under which there is no homeomorphism between spaces $C_{p}(M)$ and $\mathcal{F}(N)_{w}$.
@@
We provide a new characterization of the Schur property of Lipschitz-free spaces, more in the spirit of the title of our paper. Namely, in Theorem 4.14 we prove that for an infinite metric space $M$ the space $\mathcal{F}(M)$ has the Schur property if and only if for every complete discrete metric space $N$ with cardinality $d(M)$ the spaces $\mathcal{F}(M)_{w}$ and $\mathcal{F}(N)_{w}$ are sequentially homeomorphic. The main tool used for proving this result is Theorem 4.9 stating that a Banach space $E$ has the Schur property if and only if the spaces $E$ and $E_{w}$ are sequentially homeomorphic. We also observe in Theorem 4.12 that, for every two Banach spaces $E$ and $F$ both having the Schur property, the spaces $E_{w}$ and $F_{w}$ are sequentially homeomorphic if and only if $d(E)=d(F)$.
Section 5 is devoted to the second main problem of this paper, motivated by the fact that for every infinite metric space $M$, the space $\operatorname{Lip}_{0}(M)$ contains a complemented copy of the Banach space $\ell_{\infty}(d(M))$ (by Hájek–Novotný’s result [37], see also Theorem 3.2), hence there is always a continuous linear surjection $T\colon\operatorname{Lip}_{0}(M)\rightarrow\ell_{2}(2^{d(M)})$ and thus a continuous linear surjection $S\colon\operatorname{Lip}_{0}(M)\rightarrow\ell_{2}$, see Corollary 5.1.
@@
Using Corollary 5.7 and the theorem of Dalet [24] stating that the space $\operatorname{Lip}_{0}(\ell_{1})$ contains a complemented copy of $\ell_{1}$, we provide in Theorem 5.10 an extensive list of examples of Banach spaces $E$ for which the spaces $\operatorname{Lip}_{0}(E)$ admit continuous operators onto $\ell_{1}$. To this list belong, e.g., the following Banach spaces:
  • $C(K)$-spaces and so $L_{\infty}(\mu)$-spaces,
  • $L_{1}(\mu)$-spaces,
@@
It seems that apart from $\operatorname{Lip}_{0}([0,1])\simeq\ell_{\infty}$ there is no known example of a Banach space $\operatorname{Lip}_{0}(M)$ which is a Grothendieck space. Recall that for the class of $C(K)$-spaces we have the following result, due to Cembranos [17] and Räbiger [57]: for every compact space $K$, the Banach space $C(K)$ is Grothendieck if and only if $C(K)$ does not contain any complemented copy of $c_{0}$. This characterization cannot work for spaces $\operatorname{Lip}_{0}(M)$, since for each metric space $M$ the space $\operatorname{Lip}_{0}(M)$ is isometrically isomorphic to the dual Banach space $\mathcal{F}(M)^{*}$, hence it does not contain any complemented copy of $c_{0}$ (see e.g. [23, Theorem 2.4.15]), even though, by Theorem 3.2, $\operatorname{Lip}_{0}(M)$ always contains some copy of $c_{0}$. It follows that to decide that a given space $\operatorname{Lip}_{0}(M)$ is not Grothendieck we have to look for continuous linear surjections from $\operatorname{Lip}_{0}(M)$ onto $c_{0}$ or prove that the dual space $\operatorname{Lip}_{0}(M)^{*}$ is not weakly sequentially complete; see the discussion at the beginning of Section 5.1 for more details.
Based on results from the beginning of Section 5, in Subsection 5.1 we provide a number of conditions for metric spaces $M$ implying that their spaces $\operatorname{Lip}_{0}(M)$ are not Grothendieck. For example, since the Grothendieck property is preserved by continuous linear surjections and the space $\operatorname{Lip}_{0}(E)$ contains a complemented copy of the dual $E^{*}$ for each Banach space $E$, the space $\operatorname{Lip}_{0}(E)$ is not Grothendieck provided that $E^{*}$ is not Grothendieck. This observation yields a class of Banach spaces $E$ for which $\operatorname{Lip}_{0}(E)$ is not Grothendieck, e.g. all non-reflexive spaces with separable duals belong to this class, see Proposition 5.14.
@@ 2. Preliminaries
Unless stated otherwise all vector spaces considered in this article are assumed to be over the field of real numbers. For a Banach space $E$, the closed unit ball of $E$ will be denoted by $B_{E}$. $E^{*}$ denotes the dual Banach space of $E$ and $E_{w}$ denotes $E$ endowed with the weak topology, i.e. the topology $w=\sigma(E,E^{*})$. The weak* topology of $E^{*}$ will be denoted by $w^{*}$, i.e. $w^{*}=\sigma(E^{*},E)$.
@@
The Banach spaces $c_{0}$ and $L_{p}(\Omega,\Sigma,\mu)$ ($1\leq p\leq\infty$), in particular $\ell_{1}(\Gamma)$ and $\ell_{\infty}(\Gamma)$ for a set $\Gamma$ and $\ell_{\infty}^{n}$ for an integer $n\in\mathbb{N}$, are defined in the standard way. Recall that a Banach space $E$ is $\ell_{p}$-saturated for some $1\leq p<\infty$ if every infinite-dimensional closed linear subspace of $E$ contains an isomorphic copy of the space $\ell_{p}$.
For a topological space $X$, by $w(x)$ we denote the weight of $X$ (i.e. the minimal cardinality of a base of the topology of $X$), by $d(X)$ the density of $X$ (i.e. the minimal cardinality of a dense subset of $X$), by $d_{seq}(X)$ the sequential density of $X$ (i.e. the minimal cardinality of a sequentially dense subset of $X$, that is, such a subset $A\subseteq X$ that for every $x\in X$ there exists a sequence $(x_{n})_{n=0}^{\infty}$ in $A$ converging to $x$), by $c(X)$ the cellularity of $X$ (i.e. the supremum of cardinalities of families of pairwise disjoint non-empty open subsets of $X$), and by $ww(X)$ the weak weight of $X$ (i.e. the minimal weight $w(Y)$ of a Tychonoff space $Y$ such that there exists a continuous bijection $\varphi\colon X\rightarrow Y$). We always have $c(X)\leq d(X)\leq w(X)$, $d(X)\leq d_{seq}(X)$, and $w(X)\geq ww(X)$. If $X$ is metric, then we have $c(X)=d(X)=d_{seq}(X)=w(X)\geq ww(X)$, and the cellularity is attained, that is, there exists a collection of pairwise disjoint non-empty open subsets of $X$ of cardinality $c(X)$ (see [40, Theorem 8.1]).
@@ 3. Density of spaces of Lipschitz functions
In this paper we will frequently use the fact that for every infinite metric space $M$ the spaces $\operatorname{Lip}(M)$ and $\operatorname{Lip}_{0}(M)$ contain isomorphic copies of the Banach space $\ell_{\infty}(d(M))$, where $d(M)$ denotes the density of $M$ (see Theorems 3.1 and 3.2). These results can be easily deduced from the proof of Hájek and Novotný’s [37, Proposition 3] when combined with Rosenthal’s [60, Corollary 1.2]. Since our proof is somewhat more direct, we add it for the convenience of the reader and to keep this paper more self-contained. The main difference between our proof and the one in [37] is that we use the cellularity of $M$ to obtain directly a suitable family of disjoint open balls, which makes the argument simpler.
@@
Since $M$ is a metric space, its cellularity $c(M)$ is equal to the density $d(M)$, and so we may pick a family of pairwise disjoint (open) balls $(B(y_{\alpha},r_{\alpha}))_{\alpha<d(M)}$ of cardinality $d(M)$ such that $r_{\alpha}\leq 1$ for every $\alpha<d(M)$ (see [40, Theorem 8.1]). For each $\alpha<d(M)$ we set
@@
where $e_{\alpha}$ denotes the $\alpha$-th standard basic vector of $\ell_{\infty}(d(M))$, we may use Rosenthal’s [36, Theorem 7.10] to conclude the existence of a set $\Gamma$ with cardinality $d(M)$ such that $\ell_{\infty}(\Gamma)$ is isomorphic to a subspace of $\operatorname{Lip}(M)$. This finishes the proof. ∎
Following [66, Definition 2.17], for a complete metric space $M$ with the metric $\rho$ and the base point $e$, the space $M^{\dagger}$ is defined as $M\setminus\{e\}$ and equipped with the metric
@@
for $x,y\in M\setminus\{e\}$. By [66, Theorem 2.20], the Banach spaces $\operatorname{Lip}_{0}(M)$ and $\operatorname{Lip}(M^{\dagger})$ are isomorphic—we will use this result together with Theorem 3.1 to deduce the following theorem of Hájek and Novotný [37], which will constitute one of our main tools in what follows.
@@
The latter inequality yields that $\rho(x_{n},x)\rightarrow 0$ as $\rho^{\dagger}(x_{n},x)\to 0$. It follows that the inverse mapping $f^{-1}$ is also sequentially continuous and hence continuous. Consequently, $f$ is a homeomorphism between $M\setminus\{e\}$ and $M^{\dagger}$, and so $d(M)=d(M^{\dagger})$. By aforementioned [66, Theorem 2.20], the Banach spaces $\operatorname{Lip}_{0}(M)$ and $\operatorname{Lip}(M^{\dagger})$ are isomorphic and hence the claim follows from Theorem 3.1. ∎
As a direct consequence of Theorem 3.2 we obtain Hájek and Novotný’s ([37, Proposition 3]).
@@
Since by Theorem 3.2 the dual space $\operatorname{Lip}_{0}(M)$ of $\mathcal{F}(M)$ contains a copy of $\ell_{\infty}(d(M))$, the result follows from [60, Corollary 1.2]. ∎
@@
We already know that $d(\operatorname{Lip}_{0}(M))\geq 2^{d(M)}$ and $d(\operatorname{Lip}(M))\geq 2^{d(M)}$. The other direction follows from the fact that there are at most $2^{d(M)}$ many continuous real-valued functions on $M$, see e.g. [40, Theorem 10.1]. ∎
Corollary 3.4, together with the classical Toruńczyk theorem ([65]) stating that two infinite-dimensional Banach spaces are homeomorphic if and only if they have the same density, yields the following result.
@@
Of course, if $d(M)=d(N)$, then $2^{d(M)}=2^{d(N)}$, however the converse implication in general does not hold, as there are models of set theory (like original Cohen’s model for violating the Continuum Hypothesis, see [51, Chapter VII]) in which the equality $2^{\aleph_{0}}=2^{\aleph_{1}}$ is satisfied.
@@
Assume that $E$ is a closed linear subspace of $\operatorname{Lip}_{0}(M)_{p}$ isomorphic to $(c_{0})_{p}$. Since the closed unit ball $B$ of $\operatorname{Lip}_{0}(M)$ is compact in $\operatorname{Lip}_{0}(M)_{p}$, for the sequence $(B_{n})_{n=0}^{\infty}$ of compact sets, where $B_{n}=nB$ for $n\in\mathbb{N}$, we obtain the compact cover $(B_{n}\cap E)_{n=0}^{\infty}$ of $E$. On the other hand, the space $(c_{0})_{p}$ is isomorphic to the space $C_{p}(\mathbb{N}^{\#})$, where $\mathbb{N}^{\#}$ is a one-point compactification of the discrete space $\mathbb{N}$ of natural numbers. It follows that $C_{p}(\mathbb{N}^{\#})$ is covered by a sequence of compact sets. This however yields a contradiction by [4, Theorem I.2.1]. ∎
@@
We will now compute densities of spaces of Lipschitz functions endowed with various topologies weaker than the compact-open topology. The next theorem is an extension of Noble’s [55, Theorem 1].
@@
Since $M$ is metric, $d(M)=w(M)\geq ww(M)$. By [55, Theorem 1], $ww(M)=d(C(M),\tau)$.
@@
We have $G^{U,V}_{n}\subseteq G^{U,V}_{n+1}\subseteq G^{U,V}$. Also, if $G^{U,V}_{n}\neq\emptyset$, then $\rho(x,G^{U,V}_{n})\geq 1/n-{1}/{(3n^{2})}$ for every $x\in H^{U,V}$.